Subtract.
Could someone help me with this on please? I am not sure how many qt. are in a gallon? Any clues? 5? 6? HELP please.
8gal. 7qt.
-3gal 2qt.
______________

Answer by SHUgrad05(58)

There are 4 quarts in a gallon.
So:
8gal. 7qt.
-3gal 2qt.
Convert everything into quarts
8gal 7qt= 8(4)+7= 39
3gal 2qt= 3(4)+2= 14
39-14= 25qt= 6gal 1qt